OpenStudy (anonymous):

If someone could please explain thanks :) A typical color television draws about 2.1 A when connected to a 156 V source. What is the effective resistance of the T.V. set? Answer in units of Ω

OpenStudy (anonymous):

Ohm's law states that V = I * R, where: V = Voltage I = Current R = Resistance Rearranging to solve for R, you get: R = V / I = 156V / 2.1A = 74.286 Ω Thus, the effective resistance is 74.286 Ω.

OpenStudy (anonymous):

Oh okay great thanks ! what does T.V mean thats what really stumped me